OpenStudy (anonymous):

John pays $650 in property tax. What is the assessed value of his property if property taxes are 1.2% of assessed value? why is the equation set up 650=(1.2%)x i know the answer to the problem i just want to know why u set the equation up like that.

jimthompson5910 (jim_thompson5910):

If x = value of the property, then 12% of Value of property = Property Tax 0.12*x = 650 You would then solve for x by dividing both sides by 0.12

OpenStudy (anonymous):

You set up the equation right, they gave you the property tax, and they tell you that property tx is 1.2% which means in order to find the initial value you must multiply :D

OpenStudy (anonymous):

I didnt set up the equation, my teacher did but idk why she did, like why she translated it in math terms to work out the problem

OpenStudy (anonymous):

multiply or divide, its the same, but since your working back words, you do what @jim_thompson5910 said

jimthompson5910 (jim_thompson5910):

you just take 12% of the value of the property, ie you multiply the property value by 0.12

jimthompson5910 (jim_thompson5910):

that will give you the property tax of $650

OpenStudy (anonymous):

YEA, but she doesnt understand why....

jimthompson5910 (jim_thompson5910):

say the property is worth $100 take 12% of that to get 12, so the property tax is $12

jimthompson5910 (jim_thompson5910):

since 0.12 times 100 = 0.12*100 = 12

OpenStudy (anonymous):

what in the problem says oh u multiply 1.2% and x and u write out ur equation "650=1.2%x"

jimthompson5910 (jim_thompson5910):

oh its 1.2% not 12%, my bad

jimthompson5910 (jim_thompson5910):

but the same idea applies

OpenStudy (anonymous):

You just kind of have to look at the problem. They already told you what the end result is after the property tax, which is 650, and they give you the tax amount, they will never tell you what to do, it just makes sense, if you dont mind me asking, what math course do you take?

jimthompson5910 (jim_thompson5910):

whenever you have something like "x percent of y", you do the following x percent of y = (x/100)*y basically, divide x by 100, then multiply that with y

OpenStudy (anonymous):

110 and im studying to take the TEAS test, its the entry exam to apply to the nursing program

OpenStudy (anonymous):

ok in the problem where is the "translate word" for the( = sign) for 650

OpenStudy (anonymous):

I mean i hope you dont take this offensively, but setting up an equation is quite simple, i recommend you go on youtube, on how to set up equations from word problems, there are a lot of videos, and they are really helpful. Do you understand how you got the equation for this question?

OpenStudy (anonymous):

for example in the problem it says "1.2% of assed value" that means multiply. but where does it translate for the 650= in the problem. i get why in the equation 1.2%x but not 650=

jimthompson5910 (jim_thompson5910):

650 is 1.2% of property value x is the property value the word "is" translates to the = sign the word "of" translates to the * sign ------------------------------------------------------- so 650 is 1.2% of property value is the same as 650 = (1.2/100) * x where x is the property value

OpenStudy (anonymous):

there isnt an is in the problem in the sentence for 650...

OpenStudy (anonymous):

just says john pays 650 in prop tax. no is

jimthompson5910 (jim_thompson5910):

but if you say "john pays $650 in tax" and the way you find the tax is that you multiply the value by the 1.2/100, then you can say the two are the same

jimthompson5910 (jim_thompson5910):

ie you can say 650 = (1.2/100)*x

OpenStudy (anonymous):

ok i think i get it